# Counting primes with a given primitive root, uniformly

### v1

@@ 6. Almost-primitive roots
In the statement and proof of Theorem 1.4, there is no need to restrict $g$ to $\mathcal{G}$; all the arguments work just as well if we average $\min\{p_{g},x^{\frac{1}{2}-\varepsilon}\}$ over all $g$, $|g|\leq x$. For this unrestricted average, the exponent $\frac{1}{2}$ in the cutoff is optimal, in that even, square values of $g$ push the average of $\min\{p_{g},x^{\frac{1}{2}+\varepsilon}\}$ to infinity. One could hope to transcend $\frac{1}{2}$ after restoring the condition $g\in\mathcal{G}$, but it is not clear how to work that $g\in\mathcal{G}$ into the proof of a result like Proposition 5.1.
Recall from the introduction that $g$ is called an almost-primitive root mod $p$ when $g$ generates a subgroup of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ of index at most $2$. Define $p_{g}^{*}$ analogously to $p_{g}$ but with “almost-primitive root” in place of “primitive root.” We then expect that $p_{g}^{*}<\infty$ for every nonzero $g\in\mathbb{Z}$. This seems difficult to establish unconditionally, but it can be seen to follow from GRH by a slight modification of Hooley’s argument.
Our final theorem is an upper bound on the frequency of large values of $p_{g}^{*}$, strong enough to imply that $\min\{p_{g}^{*},x^{1-\varepsilon}\}$ has its expected mean value.
@@
We turn now to the proof of Theorem 6.1. This requires a new ingredient, Gallagher’s “larger sieve” (see [6] or [4, §9.7]).
@@
Let $N\in\mathbb{N}$, and let $\mathcal{S}$ be a finite set of prime powers. Suppose that all but $\overline{\nu}(q)$ residue classes mod $q$ are removed for each $q\in\mathcal{S}$. Then among any $N$ consecutive integers, the number remaining unsieved does not exceed
$\left({\sum_{q\in\mathcal{S}}\Lambda(q)-\log{N}}\right)\Bigg{/}\left({\sum_{q\in\mathcal{S}}\frac{\Lambda(q)}{\overline{\nu}(q)}-\log{N}}\right),$ (27)
@@
#$\#\{p\leq Y:P^{-}(\frac{p-1}{2})>Y^{\theta}\}\gg Y/\log^{2}{Y}.$
(The implied constant here is allowed to depend on $\theta$.) As remarked in the proof of Proposition 5.1, the Bombieri–Vinogradov theorem in conjunction with the linear sieve implies that any $\theta<\frac{1}{4}$ is admissible. It is known how to do a little better; for instance, [4, Theorem 25.11] shows that $\theta=3/11$ is admissible. Any admissible $\theta>0$ would yield a version of Theorem 6.1; to obtain the clean exponents of $3$ and $4$, we will use the existence of an admissible $\theta>\frac{1}{4}$.
@@
Fix an admissible $\theta>\frac{1}{4}$. We let $x$ be large, and we sieve the $N:=2\lfloor x\rfloor+1$ integers in the interval $[-x,x]$. With $y:=\log^{4}{x}$, let
$\mathcal{S}=\left\{\text{primes $p$}:3<p\leq y,P^{-}\left(\frac{p-1}{2}\right)>y^{\theta}\right\},$
so that
#$\#\mathcal{S}\gg\frac{y}{\log^{2}{y}}.$
(Here $P^{-}(\cdot)$ denotes the smallest prime factor.) For each $p\in\mathcal{S}$, we remove every residue class except $0\bmod{p}$ and the classes corresponding to integers whose order mod $p$ does not exceed
@@
$\overline{\nu}(p)=1+\sum_{\begin{subarray}{c}f\mid p-1\\ f\leq z\end{subarray}}\varphi(f).$ (28)
Suppose the integer $g$, $|g|\leq x$, is removed in the sieve. Then there is a prime $p\in\mathcal{S}$ not dividing $g$ for which the order $\ell$ (say) of $g$ mod $p$ exceeds $z$. Then $\frac{p-1}{\ell}<y/z=y^{\theta}$, while every odd divisor of $p-1$ exceeds $y^{\theta}$. Thus (keeping in mind that $p\equiv 3\pmod{4}$), $\ell=\frac{p-1}{2}$ or $p-1$, meaning that $g$ is an almost-primitive root mod $p$. In particular, $p_{g}^{\ast}\leq y$.
Hence, the number of $g$, $|g|\leq x$, with $p_{g}^{\ast}>y$ is bounded above by the count of unsieved integers, which can be approached with the larger sieve. The arguments below draw inspiration from Gallagher’s proof of Theorem 2 in [6].
@@
$\Bigg{(}\sum_{p\in\mathcal{S}}\frac{\log{p}}{\overline{\nu}(p)}\Bigg{)}\Bigg{(}\sum_{p\in\mathcal{S}}\overline{\nu}(p)\log{p}\Bigg{)}\geq\Bigg{(}\sum_{p\in\mathcal{S}}\log{p}\Bigg{)}^{2}\gg(\log(y)\cdot\#\mathcal{S})^{2}\gg y^{2}/\log^{2}{y}.$
(We use here that $\log{p}\gg\log{y}$ for each $p\in\mathcal{S}$, which follows from $P^{-}(\frac{p-1}{2})>y^{\theta}$.) On the other hand, referring back to (28),
$\sum_{p\in\mathcal{S}}\overline{\nu}(p)\log{p}$ $\leq\sum_{p\in\mathcal{S}}\log{p}+\sum_{f\leq z}\varphi(f)\sum_{\begin{subarray}{c}p\in\mathcal{S}\\ p\equiv 1\pmod{f}\end{subarray}}\log{p}$
$\ll(\log{y})\#\mathcal{S}+\log{y}\sum_{f\leq z}\varphi(f)\#\{p\in\mathcal{S}:p\equiv 1\pmod{f}\}.$
Brun’s sieve implies that #$\#\mathcal{S}\ll y/\log^{2}{y}$. Brun’s sieve also handles the counts in appearing in the sum on $f$: If $p\in\mathcal{S}$, $p\equiv 1\pmod{f}$, and $p>y^{\theta}$, then $t:=\frac{p-1}{f}<y/f$, and both $tf+1,t$ have no odd prime factors up to $y^{\theta}$. The sieve shows that the number of such $t$ is
@@
#$\#\{p\in\mathcal{S}:p\equiv 1\pmod{f}\}\ll\frac{y}{\varphi(f)\log^{2}{y}},$
@@
$\log{y}\sum_{f\leq z}\varphi(f)\#\{p\in\mathcal{S}:p\equiv 1\pmod{f}\}\ll yz/\log{y}.$
@@
$\sum_{p\in\mathcal{S}}\overline{\nu}(p)\log{p}\ll yz/\log{y},$
and hence
$\sum_{p\in\mathcal{S}}\frac{\log{p}}{\overline{\nu}(p)}\gg\frac{y^{2}/\log^{2}{y}}{yz/\log{y}}=\frac{y}{z}\frac{1}{\log{y}}=\frac{y^{\theta}}{\log{y}}.$
Since $y=\log^{4}{x}$ and $\theta>\frac{1}{4}$, this last expression is of larger order than $\log{N}$, and the denominator in (27) is $\gg y^{\theta}/\log{y}$. The numerator in (27) is $\ll(\log{y})\#\mathcal{S}\ll y/\log{y}$. Therefore, the number of unsieved $g$, $|g|\leq x$, is
@@
By our choices of $y$ and $\theta$, this last expression is $o(\log^{3}{x})$. ∎
@@
Fix an admissible $\theta\in(0,1)$, and set $y=((\log{x})(\log\log{x})^{2})^{1/\theta}$. A slight tweak to the above argument shows that there are $O(y^{1-\theta})$ integers $g$, $|g|\leq x$, with $p_{g}^{\ast}>y$. Taking $\theta=3/11$, we can replace the exponents $3$ and $4$ in Theorem 6.1 with $8/3+o(1)$ and $11/3+o(1)$, respectively. It is probably the case that every $\theta\in(0,1)$ is admissible; if so, those exponents can be brought arbitrarily close to 00 and $1$.
@@ Acknowledgments
P.P. is supported by NSF award DMS-2001581. The authors thank Igor Shparlinski for a stimulating email message that prompted these investigations. A reply of Andrew Granville on MathOverflow was also useful [9].

### v2

@@ 7. Almost-primitive roots
Recall from the introduction that $g$ is called an almost-primitive root mod $p$ when $g$ generates a subgroup of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ of index at most $2$. Define $p_{g}^{*}$ analogously to $p_{g}$ but with “almost-primitive root” in place of “primitive root.” We then expect that
$p_{g}^{*}<\infty~\,\text{for \emph{every} nonzero $g\in\mathbb{Z}$}.$ (7.1)
This seems difficult to establish unconditionally, but it can be seen to follow from GRH by a modification of Hooley’s argument. As we are not aware of a reference, we include a short GRH-conditional proof of (7.1) at the end of this section.
Our final main result is an upper bound on the frequency of large values of $p_{g}^{*}$.
@@
We turn now to the proof of Theorem 7.1. This requires a new ingredient, Gallagher’s “larger sieve” (see [7] or [5, §9.7]).
@@
Let $N\in\mathbb{N}$, and let $\mathcal{D}$ be a finite set of prime powers. Suppose that all but $\overline{\nu}(d)$ residue classes mod $d$ are removed for each $d\in\mathcal{D}$. Then among any $N$ consecutive integers, the number remaining unsieved does not exceed
$\left({\sum_{d\in\mathcal{D}}\Lambda(d)-\log{N}}\right)\Bigg{/}\left({\sum_{d\in\mathcal{D}}\frac{\Lambda(d)}{\overline{\nu}(d)}-\log{N}}\right),$ (7.2)
@@
$\left\{p\leq Y:P^{-}\left(\frac{p-1}{2}\right)>Y^{\theta}\right\}\gg\frac{Y}{\log^{2}{Y}}.$
(The implied constant here is allowed to depend on $\theta$.) As remarked in the proof of Proposition 6.1, the Bombieri–Vinogradov theorem in conjunction with the linear sieve implies that any $\theta<\frac{1}{4}$ is admissible. It is known that there are admissible values of $\theta>\frac{1}{4}$; for instance, [5, Theorem 25.11] shows that $\theta=\frac{3}{11}$ is admissible.
@@
We prove a somewhat more general result. Fix an admissible $\theta\in(0,1)$. Let $x$ be a large real number, and define
$y=((\log{x})(\log\log{x})^{2})^{1/\theta}.$ (7.3)
We show that
$\{g:|g|\leq x,p_{g}^{\ast}>y\}\ll y^{1-\theta}.$ (7.4)
Theorem 7.1 follows from (7.4) upon choosing an admissible $\theta>\frac{1}{4}$.
We sieve the $N:=2\lfloor x\rfloor+1$ integers in the interval [x,x][-x,x]. Let
$\mathcal{D}=\left\{\text{primes $p$}:3<p\leq y,P^{-}\left(\frac{p-1}{2}\right)>y^{\theta}\right\}.$
Since $\theta$ is admissible,
$\mathcal{D}\gg\frac{y}{\log^{2}{y}}.$
For each $p\in\mathcal{D}$, we remove every residue class except $0\bmod{p}$ and the classes corresponding to integers whose multiplicative order mod $p$ does not exceed
@@
$\overline{\nu}(p)=1+\sum_{\begin{subarray}{c}f\mid p-1\\ f\leq z\end{subarray}}\varphi(f).$ (7.5)
Suppose the integer $g$, $|g|\leq x$, is removed in the sieve. In this case, there is a prime $p\in\mathcal{D}$ not dividing $g$ for which the order of $g$ mod $p$, which we will call $e$, exceeds $z$. Then
$\frac{p-1}{e}<\frac{y}{e}<\frac{y}{z}=y^{\theta}.$
Since every odd prime divisor of p1p-1 exceeds $y^{\theta}$, the ratio $\frac{p-1}{e}$ cannot be divisible by any odd prime. Thus, $\frac{p-1}{e}=2^{j}$ for a nonnegative integer $j$. Since $2^{j}\mid p-1$ and $p\equiv 3\pmod{4}$, either $j=0$ or $j=1$. That is, $e=\frac{p-1}{2}$ or p1p-1. Hence, $g$ is an almost-primitive root mod $p$. In particular, $p_{g}^{\ast}\leq p\leq y$.
Therefore, the number of $g$, $|g|\leq x$, with $p_{g}^{\ast}>y$ is bounded above by the count of unsieved integers, which can be approached with the larger sieve. The arguments below draw inspiration from Gallagher’s proof of [7, Theorem 2].
@@
$\Bigg{(}\sum_{p\in\mathcal{D}}\frac{\log{p}}{\overline{\nu}(p)}\Bigg{)}\Bigg{(}\sum_{p\in\mathcal{D}}\overline{\nu}(p)\log{p}\Bigg{)}\geq\Bigg{(}\sum_{p\in\mathcal{D}}\log{p}\Bigg{)}^{2}\gg((\log y)\#\mathcal{D})^{2}\gg\frac{y^{2}}{\log^{2}{y}}.$ (7.6)
(We use here that $\log{p}\gg\log{y}$ for each $p\in\mathcal{D}$, which follows from $P^{-}(\frac{p-1}{2})>y^{\theta}$.) On the other hand, referring back to (7.5),
$\sum_{p\in\mathcal{D}}\overline{\nu}(p)\log{p}$ $\leq\sum_{p\in\mathcal{D}}\log{p}+\sum_{f\leq z}\varphi(f)\sum_{\begin{subarray}{c}p\in\mathcal{D}\\ p\equiv 1\pmod{*}{f}\end{subarray}}\log{p}$
$\ll(\log{y})\#\mathcal{D}+\log{y}\sum_{f\leq z}\varphi(f)\#\{p\in\mathcal{D}:p\equiv 1\pmod{*}{f}\}.$
Brun’s sieve implies that $\mathcal{D}\ll y/\log^{2}{y}$. Brun’s sieve also handles the counts appearing in the sum on $f$: If $p\in\mathcal{D}$, $p\equiv 1\pmod{f}$, and $p>y^{\theta}$, then $t:=\frac{p-1}{f}<y/f$, and both $tf+1$ and $t$ have no odd prime factors up to $y^{\theta}$. Brun’s sieve shows that the number of such $t$ is
@@
$\{p\in\mathcal{D}:p\equiv 1\pmod{*}{f}\}\ll\frac{y}{\varphi(f)\log^{2}{y}},$
@@
$\log{y}\sum_{f\leq z}\varphi(f)\#\{p\in\mathcal{D}:p\equiv 1\pmod{*}{f}\}\ll\frac{yz}{\log{y}}.$
@@
$\sum_{p\in\mathcal{D}}\overline{\nu}(p)\log{p}\ll\frac{yz}{\log{y}},$
and hence by (7.6),
$\sum_{p\in\mathcal{D}}\frac{\log{p}}{\overline{\nu}(p)}\gg\frac{y^{2}/\log^{2}{y}}{yz/\log{y}}=\frac{y^{\theta}}{\log{y}}.$
Recalling our definition (7.3) of $y$, we have that
$\frac{y^{\theta}}{\log{y}}\gg(\log{x})(\log\log{x}),$
which is of larger order than $\log{N}$. Hence, the denominator in (7.2) is $\gg y^{\theta}/\log{y}$. The numerator in (7.2) is bounded above by $\sum_{p\in\mathcal{D}}\log{p}\leq(\log{y})\#\mathcal{D}\ll y/\log{y}$. Therefore, the number of unsieved $g$, $|g|\leq x$, is
@@
This completes the proof of (7.4). ∎
@@
It seems likely that every $\theta\in(0,1)$ is admissible. If so, (7.4) implies that the exponents $3$ and $4$ in Theorem 7.1 can be brought arbitrarily close to 0 and $1$, respectively.
Proof of (7.1), assuming GRH.
Fix $g\in\mathbb{Z}$, $g\neq 0$. If $g\in\{\pm 1\}$, then $p_{g}^{\ast}=2$. So we may assume that $|g|>1$. As before, we let $h$ denote the largest positive integer for which $g\in(\mathbb{Q}^{\times})^{h}$. Since $g$ is fixed, we will allow implied constants below to depend on $g$ (and hence also on $h$, as fixing $g$ fixes $h$).
To prove $p_{g}^{\ast}$ exists, it is enough to show there is some prime $p\equiv 3\pmod{4}$, $p\nmid g$, with the property that $p$ passes the $\ell$-test for every odd prime $\ell$. Indeed, if $p$ is any such prime and $t$ is the order of $g$ mod $p$, then $\frac{p-1}{t}$ is a divisor of p1p-1 not divisible by any odd prime $\ell$. Hence, $\frac{p-1}{t}$ is a power of $2$. As $p\equiv 3\pmod{4}$, we must have $\frac{p-1}{t}=1$ or $2$, so that $t=\frac{p-1}{2}$ or t=p1t=p-1. Therefore, $g$ is an almost primitive root mod $p$. (We encountered a similar argument in the proof of Theorem 7.1.)
Let $x$ be large, and let $\mathcal{P}=\{\text{primes $3<p\leq x$}:P^{-}(\frac{p-1}{2})>x^{1/5}\}$. Then $\mathcal{P}\gg x/(\log{x})^{2}$. We will show that as $x\to\infty$, all but $o(x/(\log{x})^{2})$ primes $p\in\mathcal{P}$ have the desired property. In particular, there is at least one such $p$.
Clearly (once $x$ is large), each $p\in\mathcal{P}$ belongs to the residue class $3$ mod $4$. Suppose now that $p\in\mathcal{P}$ but that $p$ fails the $\ell$-test for an odd prime $\ell$. As $\ell\mid p-1$, we have $x^{1/5}<\ell\leq x$.
Suppose to start with that $\ell\in(x^{1/5},x^{1/2}/(\log{x})^{3}]$. The number of $p\in\mathcal{P}$ failing the $\ell$-test is certainly no more than the total number of primes $p\leq x$ failing the $\ell$-test, which can be bounded by Lemma 3.1. Indeed, using that $\mathbb{Q}(\zeta_{\ell},\sqrt[\ell]{g}):\mathbb{Q}]=(\ell-1)\frac{\ell}{(\ell,h)}\gg\ell^{2}$, we get from Lemma 3.1 that there are $O(\frac{x}{\ell^{2}\log{x}}+x^{1/2}\log{x})$ such $p$. Summing on $\ell$, the number of $p\in\mathcal{P}$ failing the $\ell$-test for some $\ell\in(x^{1/5},x^{1/2}/(\log{x})^{3}]$ is $O(x/(\log{x})^{3})=o(x/(\log{x})^{2})$.
Suppose next that $p\in\mathcal{P}$ fails the $\ell$-test for a prime $\ell\in(x^{1/2}/(\log{x})^{3},x^{1/2}(\log{x})^{3}]$. Write $p=1+2\ell m$. Then $m\leq x/2\ell$ and $m$ avoids the residue classes $0\bmod{r}$ and $(2\ell)^{-1}\bmod{r}$ for each odd prime $r\leq x^{1/5}$; furthermore, these two residue classes are distinct for each $r$. Noting that $x^{1/5}\leq x/2\ell$, Brun’s sieve bounds the number of possibilities for $m$ given $\ell$ (and hence for $p$ given $\ell$) as
$\ll\frac{x}{2\ell}\prod_{2<r\leq x^{1/5}}\left(1-\frac{2}{r}\right)\ll\frac{x}{\ell(\log{x})^{2}}.$
Summing on $\ell$ with Mertens’ theorem, we conclude that the number of $p\in\mathcal{P}$ failing the $\ell$-test for some $\ell\in(x^{1/2}/(\log{x})^{3},x^{1/2}(\log{x})^{3}]$ is $O(x\log\log{x}/(\log{x})^{3})=o(x/(\log{x})^{2})$.
Finally, we suppose that $p\in\mathcal{P}$ fails the $\ell$-test for an $\ell>x^{1/2}(\log{x})^{3}$. Then the order of $g$ mod $p$ is at most $x^{1/2}/(\log{x})^{3}$. Hence, $p\mid g^{m}-1$ for a natural number $m<x^{1/2}/(\log{x})^{3}$. For each $m\in\mathbb{N}$, the number of prime divisors of $g^{m}-1$ is $O(m)$. Summing on $m<x^{1/2}/(\log{x})^{3}$, we see that the number of $p$ arising in this way is $O(x/(\log{x})^{6})$, which is certainly $o(x/(\log{x})^{2})$. This completes the proof. ∎
@@ Acknowledgments
We thank Igor Shparlinski for a stimulating email message that prompted these investigations. We also found useful a reply left by Andrew Granville on MathOverflow [11]. The referee caught many typos in our initial submission, and we are grateful to them for their care and attention. This paper was written while $P.$ was supported by NSF award DMS-2001581.